OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

you use the formula point slop form y - y1 = m ( x- x1 ) make sure x and y are consistent you plug them in accordingly

OpenStudy (e.mccormick):

You need to find the slope of the first line. Then you use that ans the point to do what mth3v4 said.

OpenStudy (e.mccormick):

The slope should be reaaaaaly obvious when it is in \(y=mx+b\) form. You just need m. In this one, the m is hiding, it is implied rather than stated.

OpenStudy (anonymous):

i dont get it

OpenStudy (e.mccormick):

\(y = x + 8\) \(y = mx + b\) What is the m in the first one?